OpenStudy (anonymous):

The rule for G^-1(x) in G(x)=3x-5/4 is G^-1(x) = A.(4x-5)/3 B.(4x+5)/3 C.(3x+5)/-4

OpenStudy (anonymous):

write G(x) = y and then solve for x like y = 3x-5/4 solve for x now

OpenStudy (anonymous):

if y = 3x-5/4 then x =???

OpenStudy (anonymous):

SO do you just change the signs?

OpenStudy (anonymous):

no the whole equation changes

OpenStudy (anonymous):

ok look y = 3x-5/4 y+3/4 = 3x can you find x now??

OpenStudy (anonymous):

ok look y = 3x-5/4 y+5/4 = 3x now find x